Rust (iron(III) oxide) is formed by the reaction

of oxygen with iron.

4Fe(s) + 302 (g)

> 2Fe2O3 (s)

a. Calculate the mass of oxygen required to

react with 10.0 g of iron.


1
Expert's answer
2022-03-15T03:02:01-0400

M(Fe)=56g/mol

n(Fe)=10/56=0.1786 mol

n(O)=0.1786x3/4=0.1338 mol

M(O)=16 g/mol

M(O2)=32 g/mol

m(O2)=32x0.1338=4.29 g